OpenStudy (anonymous):

Algebra 2 Help Please: Part 1: Change 3x + 2y = 10 to slope-intercept form. You must show all work. Part 2: State the slope and y-intercept

OpenStudy (anonymous):

we want it in y=mx+b where b is the y intercept and m is the slope.... so just solve for y.... we get y=-3/2x+5 is what i got

OpenStudy (anonymous):

so the slope is =3/2 and the y intercept is (0.5)

OpenStudy (anonymous):

slope is -3/2 sorry